# STEP: A Unified Spiking Transformer Evaluation Platform for Fair and Reproducible Benchmarking

### v1

@@ B.2 Visualization on ImageNet-1k
To further interpret the temporal dynamics of Spiking Transformers, we employ Grad-CAM++ [57] to visualize the attention maps across four simulation steps (T=1 to T=4) for both Spikformer [34] and QKFormer [10] on ImageNet-1k samples (Fig. 7). These visualizations offer insight into how each model accumulates temporal evidence and localizes discriminative features over time.
@@ B.3 Result on COCO
To assess the impact of pretraining on detection performance, we adopt SDTv2 [13] as the backbone for object detection, integrating it into the Mask R-CNN framework. SDTv2 replaces the standard CNN backbone with a custom Spiking Vision Transformer, featuring embedding dimensions of [128, 256, 512, 640], 8 heads, and 8 layers per stage. Its spike-driven self-attention (SDSA) enables efficient feature extraction with reduced computational cost.

### v2

@@ B.2 Visualization on ImageNet-1k
To further interpret the temporal dynamics of Spiking Transformers, we employ Grad-CAM++ [chattopadhay2018grad] to visualize the attention maps across four simulation steps (T=1 to T=4) for both Spikformer [zhou2024spikformer] and QKFormer [zhou2024qkformer] on ImageNet-1k samples (Fig. 7). These visualizations offer insight into how each model accumulates temporal evidence and localizes discriminative features over time.
@@ B.3 Result on COCO
To assess the impact of pretraining on detection performance, we adopt SDTv2 [yao2024spike] as the backbone for object detection, integrating it into the Mask R-CNN framework. SDTv2 replaces the standard CNN backbone with a custom Spiking Vision Transformer, featuring embedding dimensions of [128, 256, 512, 640], 8 heads, and 8 layers per stage. Its spike-driven self-attention (SDSA) enables efficient feature extraction with reduced computational cost.
@@
B.4 Framework Robustness
To demonstrate the robustness of the framework, we subsequently tested the main models on the primary datasets. The results in Tab. 12 show that the models reproduced based on our framework maintained both standard deviation and confidence intervals within a reasonable range across multiple datasets, including both static and neuromorphic ones.
Table 12: Classification accuracy and confidence intervals on CIFAR10, CIFAR100, and CIFAR10-DVS datasets.
Model Dataset Acc@1 (%) Std (%) t-95% Confidence Interval
Spikformer CIFAR10 95.16 0.09 [95.04%, 95.27%]
SDT CIFAR10 95.77 0.04 [95.72%, 95.82%]
QKFormer CIFAR10 96.21 0.03 [96.18%, 96.25%]
Spikformer+SEMM CIFAR10 95.55 0.33 [95.14%, 95.96%]
Spikingformer CIFAR10 95.47 0.11 [95.32%, 95.61%]
Spikformer CIFAR100 77.76 0.31 [77.38%, 78.14%]
SDT CIFAR100 78.37 0.14 [78.19%, 78.54%]
QKFormer CIFAR100 79.95 0.18 [79.72%, 80.17%]
Spikformer+SEMM CIFAR100 78.41 0.52 [77.76%, 79.06%]
Spikingformer CIFAR100 79.33 0.21 [79.06%, 79.59%]
Spikformer CIFAR10-DVS 83.40 1.76 [81.12%, 85.49%]
SDT CIFAR10-DVS 80.33 0.60 [79.58%, 81.08%]
QKFormer CIFAR10-DVS 79.77 0.58 [79.05%, 80.49%]
B.5 Complicated Datasets
In addition to traditional tasks such as classification, segmentation, and detection, the capabilities of the Spiking Transformer have also been generalized to other datasets. These datasets include 3D point cloud classification tasks such as ModelNet10/40, as well as event-based video detection using DVS data. However, processing these datasets requires model-specific optimization, meaning that some basic baselines cannot be quickly adapted to support these tasks. Nevertheless, our framework integrates them to facilitate development and research for users. To demonstrate the successful integration of these tasks within our framework, we obtained preliminary results using the corresponding models on their respective tasks:
model mAP@[0.50:0.95] mAP@0.5 AR@[0.50:0.95] (all) AR@[0.50:0.95] (large)
SDT [yao2023spike] 0.000 0.001 0.008 0.028
SODFormer [li2023sodformer] 0.000 0.001 0.023 0.034
Table 13: SpikeDriven Transformer & SODFormer(baseline) on PKU_DAVIS_SOD with 3 epochs
Model Dataset Step Acc@1(%) Epoch
Spiking Point Transformer [wu2025spiking] ModelNet10 4 90.5 200
Table 14: Performance of Spiking Point Transformer on the ModelNet10 dataset
According to Tab. 13, SDT exhibits a significant performance gap compared to the baseline on tasks without specific adaptation. In addition, it should be noted that the results of SODFormer and Spiking Point Transformer are presented merely to demonstrate the framework’s compatibility with different datasets and tasks; their parameters are not fully aligned with those in the original papers and therefore do not reflect the actual performance of the models.